OpenStudy (anonymous):

Please help!!!! 2√6 + 3√96

OpenStudy (anonymous):

\[14\sqrt{6}\]

OpenStudy (anonymous):

@DustyJ how did u do it?

OpenStudy (anonymous):

Break down \[\sqrt{96}\] into \[(\sqrt{16})*(\sqrt{6})\] and the square root of 16 becomes 4 which you multiply by the the 3 in front of it. That gives you:\[(2\sqrt{6})+(12\sqrt{6})\] Then treat the square root of 6 almost like a variable and combine the 2 and 12 and tag the square root of 6 at the end giving you\[14\sqrt{6}\]

OpenStudy (anonymous):

@DustyJ thanks